OpenStudy (anonymous):

4*(32 - 2) + 12 / 4 x 3

OpenStudy (pfenn1):

Is the expression\[4(32 - 2) + 12 \over 4 x 3\]or \[4 \times (32 - 2) + \frac{12}{4 \times 3}\]?

OpenStudy (pfenn1):

*4 (3)